OpenStudy (anonymous):

A sphere has a surface area of 160 square centimeters. Which choice is the closest to its diameter? Use 3.14 to approximate pi. 2.01 centimeters 6.32 centimeters 7.14 centimeters 12.74 centimeters

OpenStudy (cwrw238):

the formula for surface area of a sphere is 4 pi r^2 so we have 160 = 4 * 3.14 * r^2 r^2 = 160 / (4 * 3.14) now calculate r ^2 and find its aquare root to get r diameter = 2 * r